OpenStudy (anonymous):

simplify (3y)^4 x y-2 a.) 1/2y^2 b.) 3y^2 c.) 1/81y^2 d.) 81y^2

OpenStudy (tkhunny):

Please, never use 'x' to mean multiplication. Very, VERY confusing. (3y)^4 = (3^4)(y^4) y^(-2) = 1/(y^2) Plenty to work with.